OpenStudy (anonymous):

cos ( x+ pi/2) = -sin x Can you help me verify

OpenStudy (anonymous):

\[\cos \left( a+b \right)=\cos a \cos b-\sin a \sin b\] \[\cos \frac{ \pi }{ 2 }=0,\sin \frac{ \pi }{ 2 }=1\]